OpenStudy (anonymous):

Does anyone want to help me on 2 half life questions?

OpenStudy (anonymous):

1. The half-life of a carbon-14 is 5730 years. Carbon-14 can be used to date objects that are a maximum of about 58,000 years old. Approximately how many half-lives is this?

OpenStudy (anonymous):

2. How much carbon-14 is present in a 58,000 year old skeleton that originally contained 456 g of carbon-14?

TheSmartOne (thesmartone):

Use this formula to calculate half-lives: \(\sf\Large A = A_0 \times 2^{\frac{-t}{h}}\) \(\sf A\) is the final amount. \(\sf A_0\) is the Initial Amount \(\sf t\) is the Time \(\sf h\) is the Half-Life

OpenStudy (anonymous):

I know the formula, but I don't know where I have to put the 456 g??

OpenStudy (anonymous):

@january123

OpenStudy (aaronq):

\(\sf \color{green}{>originally~ contained~ 456~ g}\) \(\sf \color{green}{>originally}\) i.e. initial amount, \(\sf A_o\)

OpenStudy (aaronq):

I also think that formula i wrong, it's supposed to be: \(\huge \sf A=A_0*\dfrac{1}{2}^{\large \dfrac{-t}{h}}\)

TheSmartOne (thesmartone):

My bad, aaronq has the correct formula (: